OpenStudy (anonymous):

How do you write this expression with a rationalized denominator? 5 * sqrt(3) / 2 - sqrt(3)

OpenStudy (anonymous):

I get as far as \[\frac{ 5\sqrt{3} }{ 2 - \sqrt{3} } * \frac{ 2 + \sqrt{3} }{ 2 + \sqrt{3} }\] and then i'm stuck

OpenStudy (isaiah.feynman):

Open study won't show the expression properly.

OpenStudy (anonymous):

ugh [5 * sqrt(3) / 2 - sqrt(3)] * [2 + sqrt(3) / 2 + sqrt(3)]

OpenStudy (anonymous):

@ganeshie8 help please?

ganeshie8 (ganeshie8):

thats a very good start, next multiply numerator separately multiply denominator separately

ganeshie8 (ganeshie8):

\(\large \frac{ 5\sqrt{3} }{ 2 - \sqrt{3} } * \color{red}{\frac{ 2 + \sqrt{3} }{ 2 + \sqrt{3} }} \) \(\large \frac{ 5\sqrt{3} \color{red}{(2+\sqrt{3})}}{ (2 - \sqrt{3})\color{red}{(2+\sqrt{3})} } \)

ganeshie8 (ganeshie8):

remember the identity (a+b)(a-b) = a^2 - b^2 ?

OpenStudy (anonymous):

yeah.... so 10 sqrt(3) + 15 / 4 - 3 4 -3 = 1 so 15 + 10 sqrt(3) is that right?

ganeshie8 (ganeshie8):

\(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you!!

ganeshie8 (ganeshie8):

np =)